OpenStudy (anonymous):

If sin x = -1/3 and x is in quadrant III. How do I find cosx, tanx, cotx, secx, and cscx? the negative sign is throwing me off.

OpenStudy (anonymous):

Steps: 1) draw a triangle 2) sinx = opposite/hyp so your hyp. = 3 then use this rule to find the last side: \[c^2=x^2+y^2\] 3) cosx = adj/hyp 4) tanx = sinx/cosx 5) cotx = 1/tanx 6) secx = 1/cosx 7) cscx = 1/sin x ^_^ give it a try

OpenStudy (anonymous):

you have y, find x :)

OpenStudy (anonymous):

is x square root of 8?

OpenStudy (anonymous):

yep ^_^ now all you have to do, is plug in the right sides in the following trigs :)

OpenStudy (anonymous):

cos=- √8/3 tan =1/√8 cot=√8 sec= -3√8/8 csc = -3 are those right? I don't think they are

OpenStudy (anonymous):

tan is wrong, tanx = sinx/cosx = (-1/sqrt(8)) sec is the inverse of cos x ^_^ and csc is the inverse of sinx :) so secx = sqrt(8) /3 ^_^

OpenStudy (anonymous):

so is my x = -1 y = √8 r = 3

OpenStudy (anonymous):

x = sqrt(8) , y = 1 and c = 3 ^_^

OpenStudy (anonymous):

is y negative 1?

OpenStudy (anonymous):

yep!

OpenStudy (anonymous):

so is cos negative √8/3 or positive....

OpenStudy (anonymous):

positive because x = positive and hyp = positive ^_^

OpenStudy (anonymous):

hmm you said it's in the third Quadrant right?

OpenStudy (anonymous):

if it is in the third quadrant then both sin x and cosx = negative and tan x = positive